A simple random sample of 80 items resulted in a sample of 49. The population standard

deviation is σ = 7 .

i. What is the standard error of the mean? (1 Mark)

ii. At 99% probability, what is the margin of error?


1
Expert's answer
2021-12-12T16:05:21-0500

i.

"SE=\sigma/\sqrt{n}=7/\sqrt{49}=1"


ii.

"MOE=Z_{0.99}\sqrt{\sigma^2/n}=2.576"